OpenStudy (anonymous):

Solve the system of equations. 9x + y = 2 –4x – y = –17 A. (–3, 29) B. (1, –7) C. (0, 0) D. (4, 1)

OpenStudy (anonymous):

i literally suck at math. so any help would be great....

OpenStudy (danjs):

you can simply add those two together, and the Y-variable will cancel out...

OpenStudy (danjs):

9x + y = 2 –4x – y = –17 + ----------------- 5x + 0y = -15

OpenStudy (danjs):

now you have one variable since 0y = 0 5x = -15 solve for the X value first from that

OpenStudy (danjs):

good?

OpenStudy (anonymous):

Yes thank you..